# home page: https://www.ais.uni-bonn.de/

### v1

@@ 5 Conclusion
This paper explored language-conditioned multitask models for robot manipulation using diffusion. We proposed several enhancements to the 3DDA architecture [3dda], contributing to the development of long-horizon VLAs capable of executing multiple tasks in sequence. Our proposed model EL3DD was evaluated on the CALVIN dataset [CALVIN] and demonstrated significant improvements over the 3DDA baseline.
The proposed modifications enhanced the model’s language capabilities through additional S-BERT [SentenceBert] embeddings, improved its visual perception using an LSeg [Lseg] backbone to generate per-pixel CLIP embeddings, and introduced latent diffusion [LatentStableDiffusion] into the 3DDA architecture. Combining these enhancements resulted in a model that achieved great performance on the CALVIN benchmark for multitask manipulation and is the new state of the art for models directly generating end-effector actions using diffusion.
Despite these advancements, there remain several challenges and limitations. Since our investigation focused exclusively on the CALVIN dataset, it remains uncertain how well the model generalizes to other datasets with different tasks. Similarly, we did not conduct real-world tests to evaluate the trained model’s performance in less controlled environments.
Building on the findings of this paper, future work should aim to diversify the datasets used for training and evaluation. Additionally, real-world deployment could offer valuable insights into the approach’s stability and robustness. The architecture could also benefit from replacing the existing relative positional encoding with a more sophisticated spatial encoding to improve the model’s scene understanding.

### v2

@@ 5 Conclusion
This paper explored language-conditioned multitask models for robot manipulation using diffusion. We proposed several enhancements to the 3DDA architecture [6], contributing to the development of long-horizon VLAs capable of executing multiple tasks in sequence. Our proposed model EL3DD was evaluated on the CALVIN dataset [11] and demonstrated significant improvements over the 3DDA baseline.
The proposed modifications enhanced the model’s language capabilities through additional S-BERT [16] embeddings, improved its visual perception using an LSeg [8] backbone to generate per-pixel CLIP embeddings, and introduced latent diffusion [17] into the 3DDA architecture. Combining these enhancements resulted in a model that achieved great performance on the CALVIN benchmark for multitask manipulation and is the new state of the art for models directly generating end-effector actions using diffusion.
Since our investigation focused exclusively on the CALVIN dataset, it remains uncertain how well the model generalizes to other datasets with different tasks. Similarly, we did not conduct real-world tests to evaluate the trained model’s performance in less controlled environments, presenting a severe limitation to this study.
Building on the findings of this paper, future work should aim to diversify the datasets used for training and evaluation. Additionally, real-world deployment could offer valuable insights into the approach’s stability and robustness. Further, replacing the positional with a spatial encoding could improve the model’s scene understanding.
@@
Acknowledgments
This work has been partially funded by the Federal Ministry of Research, Technology and Space of Germany (BMFTR) in the projects ”Learn2Grasp,” grant 16IS21080, and ”Robotics Institute Germany,” grant 16ME0999.